how many quarts of pure antifreeze must be added to 6 quarts of a 30% antifreeze solution to obtain a 60% antifreeze solution?
Answer by TimothyLamb(4379)

x = quarts of pure antifreeze
---
0.30*6 + 1.00x = 0.60*(x + 6)
0.30*6 + x = 0.60x + 0.60*6
x - 0.60x = 0.60*6 - 0.30*6
0.40x = 0.60*6 - 0.30*6
x = (0.60*6 - 0.30*6)/0.40
x = 4.5
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---